Find all real solutions of the equation. (Enter your answers as a comma-separated list.
x3 − 7x2 + 6x − 47 = x2 + 1

Answer by lwsshak3(11628)

Find all real solutions of the equation.
x3 − 7x2 + 6x − 47 = x2 + 1
x^3-8x^2+6x-48=0
using Rational Roots Theorem:
...0...|.....1......-8........6........-48..........
...1...|.....1......-7......-1........-49..
...2...|.....1......-6......-6........-60...
...7...|.....1......-1......-1........-55.
...8...|.....1........0.......6...........0 (8 is a root)
f(x)=(x-8)(x^2+6)
roots are 8 and 2 imaginary roots
note: I skipped trying some positive numbers because I knew the answer is 8 from my graphic calculator. Anyway, I showed you the method you can use.